OpenStudy (anonymous):

HELP! I WILL GIVE METAL! Quadrilateral JKLM has side lengths of JK = 2y + 3, KL = 8x – 10, LM = 5y – 6, and JM = 3x + 20. Find the values of x and y that make quadrilateral JKLM a parallelogram.

OpenStudy (anonymous):

The Answers: A. x = 6, y = 6 B. x = 6, y = 3 C. x = 13, y = 3 D. x = 2, y = 1

OpenStudy (shinalcantara):



OpenStudy (shinalcantara):

Since JK = LM then, 2y + 3 =5y - 6 2y - 5y = -6 - 3 - 3y = -9 y = 3 Since KL = JM then, 8x - 10 = 3x + 20 8x - 3x = 20 + 10 5x = 30 x = 6

OpenStudy (anonymous):

THANK YOU SO MUCH :)

OpenStudy (shinalcantara):

yw